OpenStudy (anonymous):

Find the slope at x^2y + y^2 = 5 at points ( 2,1) I keep getting 1/3. I don't think that is the right answer though..

OpenStudy (helder_edwin):

u mean the slope of the tangent line?

OpenStudy (anonymous):

Idk, it just says find the slope of that equation. I think you have to use implicit differentiation to find the answer

OpenStudy (helder_edwin):

yes u have to differentiate implicitly: \[\large 2xy+x^2y'+2yy'=0 \] \[\large -2xy=y'(x^2+2y) \] \[\large y'=\frac{-2xy}{x^2+2y} \] did u get the same?

OpenStudy (anonymous):

umm no.. I got -2/ 2x + 2

OpenStudy (helder_edwin):

how come?

OpenStudy (anonymous):

\[( 2x dy/dx + 2y) + 2ydy/dx = 0 \]

OpenStudy (anonymous):

then I just factor out dy/dx

OpenStudy (helder_edwin):

NO. the first term is a produc \[\large (uv)'=u'v+uv' \]

OpenStudy (anonymous):

Yea I know. So I did f = 2x g = y f' = 2 g' = dy/dx Then I multiplied ( f.g') + (g.f')

OpenStudy (helder_edwin):

but you wrote \(x^2y\) originally

OpenStudy (anonymous):

yea, but when you take the derivative of that won't it gives you 2xydy/dx? Then you would take that and put it into the product rule? I don't even know

OpenStudy (helder_edwin):

yes, but when u differentiate \(x^2\) u DO NOT differentiante \(y\) as well

OpenStudy (anonymous):

So can you show me how to take the derivative of that equation ? and put it in product rule?

OpenStudy (helder_edwin):

ok

OpenStudy (anonymous):

Thanks!

OpenStudy (helder_edwin):

\[\large x^2y+y^2=5 \] differentiating implicitly (using the chaing rule for \(y^2\)) \[\large (x^2y)'+(y^2)'=(5)' \] \[\large (x^2)'(y)+(x^2)(y)'+2yy'=0 \] \[\large 2xy+x^2y'+2yy'=0 \] solving for \(y'\): \[\large y'=\frac{-2xy}{x^2+2y} \]

OpenStudy (anonymous):

Thank you so much. I got -2/3 as the answer.

OpenStudy (helder_edwin):

yes, -2/3 is right. and u r welcome